15. Among beginning statistics students in one university, 46 students out of (a) About what percentage had family incomes in the range $40,000 446 reported family incomes ranging from $40,000 to $50,000 a year. (b) Guess the percentage that had family incomes in the range $45,000 to (e) Guess the percentage that had family incomes in the range $46,000 to (d) Guess the percentage that had family incomes in the range $47,000 to 50,000 a year? $46,000 a year. $47,000 a year $49,000 a year
Answer #1

a) Total number of students (beginning statistics students) who are reported family incomes ranging from $40000 to $50000 a year = 46

And total number of students = 446

Therefore required percentage = (46/446 ) * 100 = 10.31%

b) Here we assume that the students are spread uniformly in the ranging from $40000 to $50000 a year.

Required percentage = {(46000-45000)/(50000-40000)}*(46/446)*100 = 1.03%

c) Similarly here we get the required % =  {(47000-46000)/(50000-40000)}*(46/446)*100 = 1.03%

d) Here we get the required % =  {(49000-47000)/(50000-40000)}*(46/446)*100 = 2.06%